# GracklePaper (title pending)

### v1

@@ 1 Introduction
Hydrodynamical simulations of many astrophysical and cosmological processes has become a standard tool in astrophysics over the last two to three decades. Indeed, many flagship cosmological (radiation-)hydrodynamical simulations have been conducted with great success over the last decade and a half alone including Illustris [[, e.g.]]Vogelsberger_2014a, Vogelsberger_2014b, Genel_2014, Sijacki_2015, Renaissance [[, e.g.]]Xu_2013, Chen_2014, OShea_2015, Smith_2018, Eagle (Schaye et al., 2015; Crain et al., 2015; Schaller et al., 2015), Obelisk (Trebitsch et al., 2021), Horizon-AGN (Dubois et al., 2014, 2015, 2016), Edge (Rey et al., 2019) and Simba (Davé et al., 2019) to name but a few. All of these hydrodynamical simulations have, in one form or another, had to deal with radiative cooling and chemistry.
@@
On smaller physical scales, radiative cooling plays a crucial role in shaping the structure and dynamics of accretion disks surrounding stars and compact objects (e.g. Blandford & Payne, 1982; Balbus & Hawley, 1991). The thermal evolution of diffuse astrophysical plasmas is largely dictated by the form of the ‘cooling curve’—which describes how the cooling rate varies with density and temperature. This cooling behavior drives thermal instabilities, leading to the formation of a multiphase interstellar medium (McKee & Ostriker, 1977; Sutherland & Dopita, 1993). It is modelling this multi-phase medium that is now often the goal of many high resolution cosmological simulations (e.g. Kannan et al., 2025).
However, chemical networks are complex and depending on the number of species involved, the number of ordinary differential equations (ODEs) that need to be solved can be significant. Moreover, the varying time scales of the different chemical reactions can result in a series of ‘stiff’ equations making efficient convergence to a well defined solution computationally expensive. When embedded within an already complex (radiation)-hydrodynamical code this extra burden can make the overall computational footprint extremely expensive. While current chemistry solvers, for example Grackle (Smith et al., 2018) and Krome (Grassi et al., 2014) are computationally optimised they nonetheless carry a significant computational cost.
The need for methods to speed up the simulation of chemical and radiative processes warrants an investigation of potential machine learning (ML) approaches. Over the past few years, neural networks have opened up several innovative approaches for solving ODEs by leveraging their ability to approximate complex functions in a mesh-free, data‐driven manner. Approaches such as Neural Differential Equations (Kidger, 2022), Physics-Informed Neural Networks (Raissi, Perdikaris & Karniadakis, 2017) and Neural Operators (Kovachki et al., 2023) promise fast and efficient approximations to solutions of differential equations, such as the chemical rate equations. In this paper, we follow the work of Branca & Pallottini (2024) and investigate Neural Operators as a model for emulating the popular chemistry solver Grackle. We explore and report on alternative architectural choices for the Neural Operator that demonstrate an improvement in both accuracy and efficiency. We also highlight some challenges that may arise, and would need to be addressed, when integrating such an emulator into a hydro-dynamic simulation code where it would be invoked regularly.
@@ 2.1 Chemistry solver
The aim of this project is to create an emulator to predict solutions to the non-equilibrium chemistry equations solved by Grackle. Grackle (Smith et al., 2017) is a stand-alone chemistry solver that can be integrated with many commonly used cosmological simulation codes [e.g. Enzo (Bryan et al., 2014), Gadget (Springel, 2005), Arepo (Springel, 2011)]. Grackle includes treatment of chemical evolution, gas heating and cooling, metallicity, dust influences, and radiation effects. Grackle has been used in many simulation efforts, including SIMBA (Davé et al., 2019), FOGGIE (Peeples et al., 2019), and AGORA (Roca-Fàbrega et al., 2020), and features several different chemical networks designed to model the evolution of different species. In this work, we focus on solving the 9-species network, with no radiation or metallicity influence. The species in this network are e+, H+, H, H-, H2, H+${}_{2}^{+}$, He, He+, and He++. Specifically, the inclusion of molecular hydrogen, H2, is crucial for cosmological simulations, as molecular hydrogen rapidly cools the gas down to temperatures close to 200 K allowing for the high densities required for the first stars in the Universe (Palla, Salpeter & Stahler, 1983; Abel et al., 1998; Bromm, Coppi & Larson, 1999, 2002; Abel, Bryan & Norman, 2002). The conditions of zero metallicity and cooling dominated by molecular hydrogen are similar to that in the early Universe, before the formation of metal-producing stars (e.g. Bromm & Larson, 2004; Yoshida et al., 2006; Bromm et al., 2009).
@@
where $n_{x}$ represents the number density of species $x$, $k_{jl}$ is the reaction rate between species $j$ and $l$, which is dependent on the temperature $T$, and $I_{j}$ the radiative rate. Since radiation is not included in this work, $I_{j}=0$ for all species. We refer the reader to Smith et al. (2017) for the exact details on the rate equations and the values of $k_{jl}(T)$ in Grackle.
To solve these equations, each equation is solved in time until a final time $t_{f}$, usually the hydrodynamical timestep provided by the simulation, has been reached. These rate equations form a coupled ODE system, for which the solution can be written as an operator $\mathcal{G}$ acting upon the timestep $t_{f}$, the species densities, and the internal energy:
@@
Grackle also performs heating and cooling operations, updating the local temperature. One additional point to note is that Grackle works directly with the internal energy of the gas instead of the temperature of the gas, as this is a more common parameter to be used by simulation codes. The conversion between temperature and internal energy follows the formula
@@
where $E$ is the internal energy (in erg/g), $k$ the Boltzmann constant, $T$ the gas temperature, $\gamma$ the adiabatic index of an ideal gas (in this work assumed to be $5/3$), $\mu$ the mean molecular weight of the gas as a fraction of hydrogen mass, and $m_{\mathrm{H}}$ is the hydrogen mass.
Since all the species evolve at the same time, Grackle updates each species using partial updates combined with subcycling. For large values of $t_{f}$ the relative change for each species can be large, which can cause errors in the integrator. The subcycling method divides $t_{f}$ into smaller intervals $\tau$ such that no large changes occur for any species during a single subcycle.
@@ 2.2 Neural Operators
One significant application of neural operators lies in solving coupled ODEs, such as the chemical rate equations. This was demonstrated in the work of Branca & Pallottini (2024) who employed a model, known as DeepONet, to emulate the Krome chemistry solver (Grassi et al., 2014). DeepONet (Lu, Jin & Karniadakis, 2019) (Deep Operator Network) is a high-level neural operator architecture consisting of two fully connected neural networks (FCNNs) whose outputs are combined to generate the final output. In our configuration, one of the FCNNs, known as the branch network, is designed to take an eleven dimensional vector as input. The components of this vector consist of the densities of the nine species $n_{i}(0)$, the overall density of the gas $n(0)$ and the internal energy $E(0)$. These determine the initial conditions and some of the parameters (namely the reaction rates) of the rate equations. The second FCNN, known as the trunk network, is designed to take a single number as input: the time, $t_{f}$, over which the initial conditions are evolved. Both FCNNs produce vectors of configurable size, and DeepONet computes its final output by taking the dot product of these vectors. A visualization of the structure as used in this work is shown in Figure 1.
The standard DeepONet architecture, just described, outputs a single number. However, since the state of the system in this context consists of ten densities and an internal energy, to learn the solution operator in Equation 2, a model that can output an eleven dimensional vector is required. There are several strategies for extending the DeepONet to multiple outputs, as proposed by Lu et al. (2022). The most straightforward is to train separate, independent DeepONets for each parameter, but this is computationally expensive. It is also possible to split the output vector of the branch network into $N$ smaller vectors and then perform the dot product between each of these vectors and the trunk output, producing $N$ outputs. This requires that the size of the branch network output is $N$ times that of the last layer of the trunk network. The first method will be referred to as the Independent strategy, and the second as SplitBranch strategy. Lu et al. (2022) propose more alternate methods which are beyond the scope of this work.
@@ 2.3 Model Variants
The fiducial model we present here has the same network architecture as proposed by Branca & Pallottini (2024). While this particular choice of architecture has been shown to be reasonably accurate, we look to conduct a parameter study to explore alternatives. To this end, we introduce three additional models besides the fiducial model, which we refer to as Wide, Deep, and WideDeep. These models are variations on the fiducial model, changing the number of layers, nodes per layer, and multi-output strategy. The Wide model is both shallower and wider than the fiducial model, the Deep model has a deeper network, while WideDeep is both deeper and wider than the fiducial model. The details of these models are shown in Table 1. For these models, the shape of the network is identical between branch and trunk networks, with one exception. The final layer of the branch net for models using the branchSplit output strategy has 11 times as many nodes as the last layer of the trunk net, since our network features 11 unique outputs (the densities for the nine species, the energy, and the overall density). The final layer of the trunk net of the Wide model has 128 nodes, meaning that the branch net has $11\cdot 128=1408$ nodes. For all the models, the branch network also features an input layer with 11 nodes, while the trunk network features an input layer with 1 node. This input layer is not shown in Table 1.
@@ 3.2 Training
To generate a training dataset, we generate random samples of inputs by uniformly sampling each parameter in logarithmic space. For each parameter $x$, the minimum and maximum values ($x_{\mathrm{min}}$ and $x_{\mathrm{max}}$) are chosen from the ranges specified in Table 2. For each input $((n_{j},E),t_{f})$, associated outputs $G((n_{j},E),t_{f})$ are computed using Grackle, forming an input-output pair, also referred to as a sample. The training set consists of $10^{7}$ of such samples, while the test set, since it is only used to validate the model, can be smaller than the training set. It therefore consists of $10^{6}$ samples. All models were initialized using the Glorot initialization (Glorot & Bengio, 2010), use the ReLU activation function, and are evolved with a learning rate of $10^{-3}$ using the Adam optimizer (Kingma & Ba, 2014) for $4\crossproduct 10^{4}$ iterations, until the fiducial model showed no major improvements. To prevent overfitting, each model uses L1 and L2 regularization. This training procedure follows the procedure laid out by Branca & Pallottini (2024). Training was conducted on two AMD EPYC 7452 32-Core CPU chips, with each model taking up to 4 days to complete training.
@@
$t$2.6 6.3 -2 6 log(yr)
Table 2: Minimum and maximum values for each parameter as found in the simulation, as well as the selected minimum and maximum values used in creating the training and test data sets, based on the values present in the cosmological simulation. The 9 species mass densities are present, as well as the overall density, the internal energy of the gas, and the external timestep.
@@ 4.1 One Zone Test
To show the stability and consistency of the trained fiducial model, we perform a test on one cell (one zone) and compute $G(n_{0},E_{0})(t_{f})$ for 1000 values of $t_{f}$ while keeping all other inputs constant. $t_{f}$ ranges from $10^{-2}$ yr to $10^{6}$ yr, linearly distributed in log space. This forms 1000 inputs for the DeepONet. The resulting number densities and temperature predicted by this test are shown in Figure 3. The solid lines corresponds to values calculated using Grackle, while the dotted lines represents the predictions made by the fiducial model. The discrepancy between Grackle and the fiducial model for all species in this example is below 0.6 dex, and does not show a clear time evolution. This cell, being typical of the overall population, has many species near equilibrium, which are not rapidly changing.
The bottom panel of Figure 3 shows another example, this time with initial conditions were chosen to show a larger change in densities. Here, it’s clear that the DeepONet prediction struggles to capture finer details in the evolution, representing rapid changes, with errors ranging as high as 1.5 dex. This corroborates a similar result found by Branca & Pallottini (2024). Cells with a rapid evolution may be rare, but their complex evolution is computationally expensive. Nonetheless, capturing their behaviour accurately will be a challenge for ML methods.
Figure 3: One zone test based on two initial cells. The solid lines show the calculated value for each species according to Grackle, while the dashed lines show the prediction made by the fiducial model. The top panel shows the evolution of a typical cell. The ML prediction and Grackle agree for the evolution of this cell up to a variation of at most 0.6 dex. The bottom panel targets a cell with a higher variance over its evolution. The disagreement between ML and Grackle in the bottom panel is greater, at 1.5 dex.

### v2

@@ 1 Introduction
Hydrodynamical simulations of many astrophysical and cosmological processes have become a standard tool in astrophysics over the last two to three decades. Indeed, many flagship cosmological (radiation-)hydrodynamical simulations have been conducted with great success over the last decade and a half alone including Illustris [[, e.g.]]Vogelsberger_2014a, Vogelsberger_2014b, Genel_2014, Sijacki_2015, Renaissance [[, e.g.]]Xu_2013, Chen_2014, OShea_2015, Smith_2018, Eagle (Schaye et al., 2015; Crain et al., 2015; Schaller et al., 2015), Obelisk (Trebitsch et al., 2021), Horizon-AGN (Dubois et al., 2014, 2015, 2016), Edge (Rey et al., 2019) and Simba (Davé et al., 2019) to name but a few. All of these hydrodynamical simulations have, in one form or another, had to deal with radiative cooling and chemistry.
@@
On smaller physical scales, radiative cooling plays a crucial role in shaping the structure and dynamics of accretion disks surrounding stars and compact objects (e.g. Blandford & Payne, 1982; Balbus & Hawley, 1991). The thermal evolution of diffuse astrophysical plasmas is largely dictated by the form of the ‘cooling curve’—which describes how the cooling rate varies with density and temperature. This cooling behaviour drives thermal instabilities, leading to the formation of a multiphase interstellar medium (McKee & Ostriker, 1977; Sutherland & Dopita, 1993). It is modelling this multi-phase medium that is now often the goal of many high resolution cosmological simulations (e.g. Kannan et al., 2025).
However, chemical networks are complex and depending on the number of species involved, the number of ordinary differential equations (ODEs) that need to be solved can be significant. Moreover, the varying time scales of the different chemical reactions can result in a series of ‘stiff’ equations, requiring computationally expensive implicit solvers to converge to a well defined solution. When embedded within an already complex (radiation)-hydrodynamical code this extra burden can make the overall computational footprint extremely expensive. While current chemistry solvers, for example Grackle (Smith et al., 2018), CHIMES (Richings, Schaye & Oppenheimer, 2014a, b), and Krome (Grassi et al., 2014) are computationally optimised they nonetheless carry a significant computational cost.
The need for methods to speed up the simulation of chemical and radiative processes warrants an investigation of potential machine learning (ML) approaches. ML approaches have quickly become relevant in the field of astrochemistry, opening several innovative approaches for solving ODEs by leveraging their ability to approximate complex functions in a mesh-free, data‐driven manner. Techniques such as Latent ODEs with autoencoders (e.g. Grassi et al., 2022; Maes et al., 2024), Neural ODEs with autoencoders (Sulzer & Buck, 2023), Neural Networks (e.g. Grassi et al., 2011; de Mijolla et al., 2019; Holdship et al., 2021; Palud et al., 2023), Physics-Informed Neural Networks (Branca & Pallottini, 2023), Neural Fields (Asensio Ramos et al., 2024), and Neural Operators (Branca & Pallottini, 2024) promise fast and efficient approximations to solutions of chemical rate equations, with a high degree of accuracy. Benchmarking tools such as CODES (Janssen, Sulzer & Buck, 2024) are starting to get off the ground, rounding out a developing ML ecosystem in the field. These codes are as of now standalone, with few attempts to integrate them into larger codes such as cosmological simulations. Therefore, it is natural to explore the potential of solving the computational cost problem with an ML method.
In this paper, we follow the work of Branca & Pallottini (2024) and investigate Neural Operators as a model for emulating the popular chemistry solver Grackle. We explore and report on alternative architectural choices for the Neural Operator that demonstrate an improvement in both accuracy and efficiency. We also highlight some challenges that may arise, and would need to be addressed, when integrating such an emulator into a hydro-dynamic simulation code where it would be invoked regularly.
@@ 2.1 Chemistry solver
The aim of this project is to create an emulator to predict solutions to the non-equilibrium chemistry equations solved by Grackle. Grackle (Smith et al., 2017) is a stand-alone chemistry solver that can be integrated with many commonly used cosmological simulation codes [e.g. Enzo (Bryan et al., 2014), Gadget (Springel, 2005), Arepo (Springel, 2011)]. Grackle includes treatment of chemical evolution, gas heating and cooling, metallicity, dust influences, and radiation effects. Grackle has been used in many simulation efforts, including SIMBA (Davé et al., 2019), FOGGIE (Peeples et al., 2019), and AGORA (Roca-Fàbrega et al., 2020), and features several different chemical networks designed to model the evolution of different species. In this work, we focus on solving the 9-species network, with no radiation or metallicity influence. The species in this network are e-, H+, H, H-, H2, H+${}_{2}^{+}$, He, He+, and He++. Specifically, the inclusion of molecular hydrogen, H2, is crucial for cosmological simulations, as molecular hydrogen rapidly cools the gas down to temperatures close to 200 K allowing for the high densities required for the first stars in the Universe (Palla, Salpeter & Stahler, 1983; Abel et al., 1998; Bromm, Coppi & Larson, 1999, 2002; Abel, Bryan & Norman, 2002). The conditions of zero metallicity and cooling dominated by molecular hydrogen are similar to that in the early Universe, before the formation of metal-producing stars (e.g. Bromm & Larson, 2004; Yoshida et al., 2006; Bromm et al., 2009).
@@
where $n_{x}$ represents the number density of species $x$, $k_{jl}$ is the reaction rate between species $j$ and $l$, which is dependent on the temperature $T$, and $I_{j}$ the radiative rate. A positive $k_{jl}$ indicates a creation term, while a negative $k_{jl}$ indicates a destruction term. As a constraint, the H+${}_{2}^{+}$ species is always assumed to be at equilibrium since the reaction timescale for this species is short enough to be decoupled from the main integration. Since radiation is not included in this work, $I_{j}=0$ for all species. We refer the reader to Smith et al. (2017) for further details on the rate equations that are included in Grackle’s nine species network, as well as the values of $k_{jl}(T)$.
To solve these equations, Grackle solves each equation in time until a final time $t_{f}$, usually the hydrodynamical time step provided by the simulation, has been reached. These rate equations form a coupled ODE system, for which the solution can be written as an operator $\mathcal{G}$ acting upon the time step $t_{f}$, the species densities, and the internal energy:
@@
Grackle also performs heating and cooling operations, updating the local temperature. One point to note is that Grackle works directly with the internal energy of the gas instead of the temperature of the gas, as this is a more common parameter to be used by simulation codes. The conversion between temperature and internal energy follows the formula
@@
where $E$ is the internal energy (in erg/g), $k$ the Boltzmann constant, $T$ the gas temperature, $\gamma$ the adiabatic index of an ideal gas (in this work assumed to be $5/3$), $\mu$ the mean molecular weight of the gas as a fraction of hydrogen mass, and $m_{\mathrm{H}}$ is the hydrogen mass. Updating the internal energy is done according to
[…]
where $\dot{e}_{\mathrm{cool}}$ represents cooling terms, and $\dot{e}_{\mathrm{heat}}$ represents heating terms. Selecting which cooling effects to take into account is highly customizable for Grackle. We use the non-equilibrium primordial heating rates for the nine species model as outlined in Smith et al. (2017). This consists of collisional excitation, collisional ionization, recombination cooling, brehmsstrahlung, Compton heating/cooling off the CMB, photoionization heating, and H2 cooling. Further details on the rates used by Grackle can be found in Smith et al. (2017).
Since all the species evolve at the same time, Grackle uses a solver based on the backwards difference formula method, with alterations through partial updates and subcycling, following Anninos et al. (1997). For large values of $t_{f}$, relative to the reaction timescale of the species, the relative change for each species can be excessive, which can cause numerical errors in the integrator. The subcycling method divides $t_{f}$ into smaller intervals $\tau$ such that no changes larger than 10 % per subcycle occur for H, e-, and the internal energy.
@@ 2.2 Neural Operators
One significant application of neural operators lies in solving coupled ODEs, such as the chemical rate equations. This was demonstrated in the work of Branca & Pallottini (2024) who employed a model, known as DeepONet, to emulate the Krome chemistry solver (Grassi et al., 2014). DeepONet (Lu, Jin & Karniadakis, 2019) (Deep Operator Network) is a high-level neural operator architecture consisting of two fully connected neural networks (FCNNs) whose outputs are combined to generate the final output. In our configuration, one of the FCNNs, known as the branch network, is designed to take an 11-dimensional vector as input. The components of this vector consist of the densities of the nine species $n_{i}(0)$, the total density of the gas $n(0)$ and the internal energy $E(0)$. These determine the initial conditions and some of the parameters (namely the reaction rates) of the rate equations. We include the total density as a sanity check. The chemistry solver does not evolve the total density, therefore the lack of evolution of this parameter may serve as a first test on convergence. The second FCNN, known as the trunk network, is designed to take a single number as input: the time, $t_{f}$, over which the initial conditions are evolved. Both FCNNs produce vectors of configurable size, and DeepONet computes its final output by taking the dot product of these vectors. A visualization of the structure as used in this work is shown in Figure 1.
The standard DeepONet architecture, just described, outputs a single number. However, since the state of the system in this context consists of ten densities and an internal energy, to learn the solution operator in Equation 2, a model that can output an 11-dimensional vector is required. There are several strategies for extending the DeepONet to multiple outputs, as proposed by Lu et al. (2022). The most straightforward is to train separate, independent DeepONets for each parameter, but this is computationally expensive. It is also possible to split the output vector of the branch network into $N$ smaller vectors and then perform the dot product between each of these vectors and the trunk output, producing $N$ outputs. This requires that the size of the branch network output is $N$ times that of the last layer of the trunk network. For this work, $N=11$, as there are 11 predicted parameters. The first method will be referred to as the Independent strategy, and the second as branchSplit strategy. Lu et al. (2022) propose more alternate methods which are beyond the scope of this work.
@@ 2.3 Model Variants
The fiducial model we present here has the same network architecture as proposed by Branca & Pallottini (2024). While this particular choice of architecture has been shown to be reasonably accurate, we look to conduct a parameter study to explore alternatives. To this end, we introduce three additional models besides the fiducial model, which we refer to as Wide, Deep, and WideDeep. These models are variations on the fiducial model, changing the number of layers, nodes per layer, and multi-output strategy. The Wide model is both shallower including fewer layers) and wider (having more nodes per layer) than the fiducial model, the Deep model has a deeper network (including more layers), while WideDeep is both deeper and wider than the fiducial model. The details of these models are shown in Table 1. The Wide model has fewer parameters than the fiducial model, to find a more lightweight model capable of emulating the chemical network, as well as determining the influence of the branchSplit output strategy. The Deep model is focused on depth, to probe whether deep learning techniques may be a requirement for future studies. Finally, the WideDeep model is meant to see if a large quantity of parameters can provide an alternative in the accuracy/speedup trade-off. For these models, the shape of the network is identical between branch and trunk networks, with one exception. The final layer of the branch net for models using the branchSplit output strategy has 11 times as many nodes as the last layer of the trunk net, since our network features 11 unique outputs (the densities for the nine species, the energy, and the total density). The final layer of the trunk net of the Wide model has 128 nodes, meaning that the branch net has $11\cdot 128=1408$ nodes. The output layers of the branch and trunk networks are combined using the dot product to generate 11 output variables. For all the models, the branch network also features an input layer with 11 nodes, while the trunk network features an input layer with 1 node. This input layer is not shown in Table 1.
@@ 3.2 Training
To generate a training dataset, we generate random samples of inputs by uniformly sampling each parameter in logarithmic space independently, sampling the entire parameter space uniformly. For each parameter $x$, the minimum and maximum values ($x_{\mathrm{min}}$ and $x_{\mathrm{max}}$) are chosen from the ranges specified in Table 2. For each input $((n_{j},E),t_{f})$, associated outputs $G((n_{j},E),t_{f})$ are computed using Grackle, forming an input-output pair, also referred to as a sample. The samples are normalized by rescaling the values between -1 and 1, using
$x_{i\mathrm{,norm}}=2\frac{x_{i}-x_{i\mathrm{,min}}}{x_{i\mathrm{,max}}-x_{i\mathrm{,min}}}-1$ (5)
following Branca & Pallottini (2024). The training set consists of $10^{7}$ of such normalized samples, while the test set, since it is only used to validate the model, can be smaller than the training set. It therefore consists of $10^{6}$ samples. We chose a training set of size $10^{7}$ to obtain a population of samples over the entire parameter space, while keeping memory and computational constraints into account. It is important to note that these samples may have non-chemical parameter combinations. That is to say, charge is not conserved and the H/He ratio is not identical in each sample. We elected to use this method to allow for more degrees of freedom for the prediction, and to capture the performance of Grackle over the entire parameter space. We discuss this point further in §4.4.
To train the models, we use the DeepXDE package, which is designed to train DeepONet models. This package is backend-agnostic, we select a Tensorflow backend. All models were initialized using the Glorot initialization (Glorot & Bengio, 2010), use the ReLU activation function, and are evolved with a learning rate of $10^{-3}$ using the Adam optimizer (Kingma & Ba, 2014) for $4\crossproduct 10^{4}$ iterations, until the fiducial model showed no major improvements on the validation set. To prevent potential overfitting, each model uses L1 and L2 regularization. This training procedure follows the procedure laid out by Branca & Pallottini (2024). Training was conducted on two AMD EPYC 7452 32-Core CPU chips, with each model taking up to 4 days to complete training. These chips were chosen due to hardware constraints.
@@
$t_{f}$2.6 6.3 -2 6 log(yr)
Table 2: Minimum and maximum values for each parameter as found in the simulation, as well as the selected minimum and maximum values used in creating the training and test data sets, based on the values present in the cosmological simulation. The 9 species mass densities are present, as well as the total density, the internal energy of the gas, and the external time step.
@@ 4.1 One Zone Test
To show the stability and consistency of the trained fiducial model, we perform a test on one cell (one zone) and compute $G(n_{0},E_{0})(t_{f})$ for 1000 values of $t_{f}$. We perform this from the same initial $G(n_{0},E_{0})$, randomly selecting a cell for the initial values. $t_{f}$ ranges from $10^{-2}$ yr to $10^{6}$ yr, linearly distributed in log space, following the ranges found for $t_{f}$ based on a single hydrodynamics time step. This forms 1000 inputs for the DeepONet. The resulting number densities and temperature predicted by this test are shown in Figure 3. The solid lines corresponds to values calculated using Grackle, while the dotted lines represents the predictions made by the fiducial model. The discrepancy between Grackle and the fiducial model for all species in this example is below 0.6 dex. There does not appear to be a clear pattern between $t_{f}$ and the error. This cell, being typical of the overall population, has many species near equilibrium, which are not rapidly changing.
The bottom panel of Figure 3 shows another example, this time with initial conditions randomly selected from the set of cells featuring $>1$ dex error for any individual species, to show a larger change in densities with highly non-equilibrium initial conditions. Here, it’s clear that the DeepONet prediction struggles to capture finer details in the evolution, representing rapid changes, with errors ranging as high as 1.5 dex. This corroborates a similar result found by Branca & Pallottini (2024). Cells with a rapid evolution may be rare, but their complex evolution is computationally expensive. Nonetheless, capturing their behaviour accurately will be a challenge for ML methods.
Figure 3: One zone test based on two initial cells. The solid lines show the calculated values for each species according to Grackle, while the dashed lines show the prediction made by the fiducial model. The top panel shows the evolution of a typical cell. The ML prediction and Grackle agree for the evolution of this cell up to a variation of at most 0.6 dex. The bottom panel targets a cell with a higher variance over its evolution. The disagreement between ML and Grackle in the bottom panel is greater, at 1.5 dex.